Could This problem be checked for me?
2x2-12x+18
-----------
2x2-18
=2(x2-6x+9)
---------
2(x2-9)
= 2(x-3)(x-3)
------------
2(x-3)(x-3)
=1
Is this answer correct in how I figured it out? Thank you for the help!

Answer by stanbon(75887)

Could This problem be checked for me?
2x2-12x+18
-----------
2x2-18
-----------------------------------------
=2(x2-6x+9)
---------
2(x2-9)
-----------------------------------------------
= 2(x-3)(x-3)
------------
2(x-3)(x+3)
-------------------------------------------------
=(x-3)/(x+3)
=================
Cheers,
Stan H.